# Positivstellensätze for polynomial matrices with universal quantifiers

### v2

@@ 4.3.1. Exponentially bounded matrix-valued moment problem
Suppose that Assumption 3.2 holds. Then for a sequence $\mathbf{S}=(S_{{\boldsymbol{\alpha}}})_{{\boldsymbol{\alpha}}\in\mathbb{N}^{n}}\subset{\mathbb{S}}^{p}$ and $C>0$, the following are equivalent:
  1. (i)
    $\mathbf{S}$ has a matrix-valued representing measure supported on $\mathcal{X}\cap[-C,C]^{n}$;
  2. (ii)
    $M(\mathbf{S})\succeq 0$, $M^{\nu}(G\mathbf{S})\succeq 0$ and there is a constant $C_{0}>0$ such that $\|S_{{\boldsymbol{\alpha}}}\|\leq C_{0}C^{|{\boldsymbol{\alpha}}|}$ for all ${\boldsymbol{\alpha}}\in\mathbb{N}^{n}$.
In case that UQs are not present, we obtain the following corollary.
Corollary 4.3.
For a sequence $\mathbf{S}=(S_{{\boldsymbol{\alpha}}})_{{\boldsymbol{\alpha}}\in\mathbb{N}^{n}}\subset{\mathbb{S}}^{p}$ and $C>0$, the following are equivalent:
  1. (i)
    $\mathbf{S}$ admits a matrix-valued representing measure supported on
    $\{{\boldsymbol{x}}\in{\mathbb{R}}^{n}\mid H({\boldsymbol{x}})\succeq 0\}\cap[-C,C]^{n},$
    where $H({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{q}$, $q\in\mathbb{N}$;
  2. (ii)
    $M(\mathbf{S})\succeq 0$, $M(H\mathbf{S})\succeq 0$ and there is a constant $C_{0}>0$ such that $\|S_{{\boldsymbol{\alpha}}}\|\leq C_{0}C^{|{\boldsymbol{\alpha}}|}$ for all ${\boldsymbol{\alpha}}\in\mathbb{N}^{n}$.
4.3.2. Lasserre-Netzer-type Positivstellensatz for polynomial matrices with UQs
Inspired by the result of Lasserre and Netzer [30], we next derive a Positivstellensatz to characterize a polynomial matrix $F({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ that is PSD on $\mathcal{X}\cap[-1,1]^{n}$ using high-degree perturbations. For $d\in\mathbb{N}$, let $\Theta_{d}\coloneqq 1+\sum_{i=1}^{n}x_{i}^{2d}$.
Let $d(G)\coloneqq\lceil\deg_{{\boldsymbol{x}}}G/2\rceil$. Given $F({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$, let us consider the following optimization problem:
@@ 4.3.2. Lasserre-Netzer-type Positivstellensatz for polynomial matrices with UQs
$r_{d,k}^{\star}\coloneqq\left\{\begin{aligned} \inf_{r,\Sigma_{0},\Sigma}&\ r\\ \text{s.t.}\,\,&\ F({\boldsymbol{x}})+r\Theta_{d}I_{p}=\Sigma_{0}+\int_{\mathcal{Y}}\left\langle\Sigma,G\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}),\\ &\ \Sigma_{0}\in{\mathbb{S}}[{\boldsymbol{x}}]_{2d}^{p},\Sigma\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}\ \text{are SOS},\\ &\ \deg_{{\boldsymbol{x}}}\Sigma\leq 2d-2d(G),\ \deg_{{\boldsymbol{y}}}\Sigma\leq 2k,\end{aligned}\right.$ (6)
and it dual problem reads as
$\sup_{\mathbf{S}}&\ -\mathscr{L}_{\mathbf{S}}(F)\\ \text{s.t.}&\ \mathbf{S}=(S_{{\boldsymbol{\alpha}}})_{{\boldsymbol{\alpha}}\in\mathbb{N}^{n}_{2d}}\subset{\mathbb{S}}^{p},\ \mathscr{L}_{\mathbf{S}}(\Theta_{d}I_{p})\leq 1,\\ &\ M_{d}(\mathbf{S})\succeq 0,\ M_{d-d(G),k}^{\nu}(G\mathbf{S})\succeq 0.\end{aligned}\right.$ (7)
As the sequence of zero matrices is feasible to (7), it holds $r^{\star}_{d,k}\geq 0$ by the weak duality.
Assumption 4.1.
There are positive number $\lambda>0$, and open and bounded subsets $\mathcal{O}_{1}$ and $\mathcal{O}_{2}$ of $\mathcal{X}$ and $\mathcal{Y}$, respectively, such that $G({\boldsymbol{x}},{\boldsymbol{y}})\succeq\lambda I_{q}$ on $\mathcal{O}_{1}\times\mathcal{O}_{2}$.
The following proposition shows that if Assumption 4.1 holds and $F({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ is PSD on $\mathcal{X}\cap[-1,1]^{n}$, then the optimal value $r^{\star}_{d,k}\to 0$ as $d,k\to\infty$. Its proof is postponed to Section 6.6.
Proposition 4.2.
Suppose that Assumptions 3.2 and 4.1 hold, and $F({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ is PSD on $\mathcal{X}\cap[-1,1]^{n}$. Then, for any $\varepsilon>0$, there exist $\ell_{1},\ell_{2}\in\mathbb{N}$ such that $r_{d,k}^{\star}\leq\varepsilon$ for all $d\geq\ell_{1},k\geq\ell_{2}$.
We can now give the Lasserre-Netzer-type Positivstellensatz for polynomial matrices with UQs.

### v4

@@ 5.3. The non-compact case
Let Assumption 3.2 hold. Then the limit $\tau_{0}(\infty,\infty)$ exists, and $\tau_{0}(\infty,\infty)\leq\tau^{\star}$. Moreover, if the optimal solution set $\mathcal{S}$ of (8) is nonempty and bounded, we have $\tau_{0}(\infty,\infty)=\tau^{\star}$.
Proof.
Since Assumption 3.2 holds, Corollary 4.1 implies that, for any $\varepsilon>0$ and sufficiently large $\ell$, $\mathcal{F}_{\varepsilon}(\ell,\ell)$ contains the feasible set $\mathcal{F}$ of (8) and hence $\tau_{\varepsilon}(\ell,\ell)\leq\tau^{\star}$. It is clear that $\tau_{\varepsilon}(\ell,\ell)$ is non-increasing as $\ell\to\infty$ for fixed $\varepsilon>0$ and is non-decreasing as $\varepsilon\to 0$ for fixed $\ell$. Hence, the limit $\tau_{0}(\infty,\infty)$ exists and $\tau_{0}(\infty,\infty)\in[-\infty,\tau^{\star}]$.
Assuming that the optimal solution set $\mathcal{S}$ of (8) is nonempty and bounded, we show that $\tau_{0}(\infty,\infty)=\tau^{\star}$. Suppose on the contrary that $\tau_{0}(\infty,\infty)<\tau^{\star}$. Then, there exists a sequence $\{\varepsilon_{j}\}_{j\in\mathbb{N}}$ with $\lim_{j\to\infty}\varepsilon_{j}=0$ and a sequence $\{{\boldsymbol{\gamma}}^{(\ell_{j})}\}_{j\in\mathbb{N}}$ of feasible points of (13) with $(\varepsilon,k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})=(\varepsilon_{j},\ell_{j},\ell_{j})$ such that for each $j\in\mathbb{N}$, ${\boldsymbol{c}}^{\intercal}{\boldsymbol{\gamma}}^{(\ell_{j})}\leq\tau^{\star}-\delta$ for some $\delta>0$. We claim that $\{{\boldsymbol{\gamma}}^{(\ell_{j})}\}_{j\in\mathbb{N}}$ is bounded. Suppose that $\|{\boldsymbol{\gamma}}^{(\ell_{j})}\|\to\infty$ as $j\to\infty$ and without loss of generality that $\lim_{j\to\infty}{\boldsymbol{\gamma}}^{(\ell_{j})}/\|{\boldsymbol{\gamma}}^{(\ell_{j})}\|={\boldsymbol{v}}$ for some ${\boldsymbol{v}}=(v_{1},\ldots,v_{r})\in{\mathbb{R}}^{r}$ with $\|{\boldsymbol{v}}\|=1$. Since ${\boldsymbol{c}}^{\intercal}{\boldsymbol{\gamma}}^{(\ell_{j})}$ is bounded by $\tau^{\star}-\delta$ for all $j\in\mathbb{N}$, we have ${\boldsymbol{c}}^{\intercal}{\boldsymbol{v}}=0$. Moreover, for arbitrarily fixed ${\boldsymbol{u}}\in\mathcal{X}$, we have
$\Lambda({\boldsymbol{u}},{\boldsymbol{\gamma}}^{(\ell_{j})})+\varepsilon_{j}\theta({\boldsymbol{u}})^{d_{\Lambda}}I_{p}=\Lambda_{0}({\boldsymbol{u}})-\sum_{i=1}^{r}\Lambda_{i}({\boldsymbol{u}})\gamma^{(\ell_{j})}_{i}+\varepsilon_{j}\theta({\boldsymbol{u}})^{d_{\Lambda}}I_{p}\succeq 0,\ \forall j\in\mathbb{N}.$ (14)
Dividing the left-hand side of (14) by $\|{\boldsymbol{\gamma}}^{(\ell_{j})}\|$ and letting $j\to\infty$, we get $\sum_{i=1}^{r}\Lambda_{i}({\boldsymbol{u}})v_{i}\succeq 0$. Since $\boldsymbol{u}$} is an arbitrary point in $\mathcal{X}$ and ${\boldsymbol{c}}^{\intercal}{\boldsymbol{v}}=0$, for any point ${\boldsymbol{\gamma}}^{\star}\in\mathcal{S}$, we have ${\boldsymbol{\gamma}}^{\star}+t{\boldsymbol{v}}\in\mathcal{S}$ for all $t>0$, which contradicts the boundedness of $\mathcal{S}$. As we have proved that $\{{\boldsymbol{\gamma}}^{(\ell_{j})}\}_{j\in\mathbb{N}}$ is bounded, we can assume without loss of generality that $\lim_{j\to\infty}{\boldsymbol{\gamma}}^{(\ell_{j})}=\hat{{\boldsymbol{\gamma}}}$ for some $\hat{{\boldsymbol{\gamma}}}\in{\mathbb{R}}^{r}$. Then letting $j\to\infty$ in (14) yields the result that $\hat{{\boldsymbol{\gamma}}}$ is feasible to (8). Thus, we have
$\tau^{\star}\leq{\boldsymbol{c}}^{\intercal}\hat{{\boldsymbol{\gamma}}}=\lim_{j\to\infty}{\boldsymbol{c}}^{\intercal}{\boldsymbol{\gamma}}^{(\ell_{j})}\leq\tau^{\star}-\delta,$
a contradiction. ∎
Now, let us consider the problem
$\tilde{\tau}^{\star}\coloneqq\inf_{{\boldsymbol{\gamma}}\in{\mathbb{R}}^{r}}\ {\boldsymbol{c}}^{\intercal}{\boldsymbol{\gamma}}\ \text{ s.t. }\ \Lambda({\boldsymbol{x}},{\boldsymbol{\gamma}})\succeq 0,\ \forall{\boldsymbol{x}}\in\mathcal{X}\cap{\mathbb{R}}_{+}^{n},$ (15)
where $\mathcal{X}$ is non-compact and defined in (8), and $\mathcal{Y}$ is defined in (7). After fixing a constant $\varepsilon>0$, we apply Corollary 4.3 to construct the following hierarchy of SDP relaxations:
@@
$\left\{\begin{aligned} \tilde{\tau}_{\varepsilon}(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})\coloneqq\inf_{{\boldsymbol{\gamma}},S_{0},S}&\ {\boldsymbol{c}}^{\intercal}{\boldsymbol{\gamma}}\\ \text{s.t.}&\ (1+{\boldsymbol{e}}^{\intercal}{\boldsymbol{x}})^{2k_{{\boldsymbol{x}}}-2d_{\Lambda}}(\Lambda({\boldsymbol{x}},{\boldsymbol{\gamma}})+\varepsilon(1+{\boldsymbol{e}}^{\intercal}{\boldsymbol{x}})^{2d_{\Lambda}}I_{p})=P_{0}({\boldsymbol{x}})+\int_{\mathcal{Y}}\left\langle P({\boldsymbol{x}},{\boldsymbol{y}}),G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}),\\ &\ P_{0}=\sum_{{\boldsymbol{\alpha}}\in\mathbb{N}^{n}_{2k_{{\boldsymbol{x}}}}}P_{0,{\boldsymbol{\alpha}}}{\boldsymbol{x}}^{{\boldsymbol{\alpha}}}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}\text{ with each }P_{0,{\boldsymbol{\alpha}}}\succeq 0,\\ &\ P=\sum_{{\boldsymbol{\alpha}}\in\mathbb{N}^{n}_{2(k_{{\boldsymbol{x}}}-d_{G})}}P_{{\boldsymbol{\alpha}}}({\boldsymbol{y}}){\boldsymbol{x}}^{{\boldsymbol{\alpha}}}\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq},\ \deg_{{\boldsymbol{y}}}P\leq 2(k_{{\boldsymbol{y}}}-\lceil\deg_{{\boldsymbol{y}}}G/2\rceil),\\ &\ \text{ with each }P_{{\boldsymbol{\alpha}}}({\boldsymbol{y}})\in\mathcal{Q}^{pq}_{k_{{\boldsymbol{y}}}-\lceil\deg_{{\boldsymbol{y}}}G/2\rceil}(\mathbf{H}).\end{aligned}\right$. (16)
For the sequence $\{\tilde{\tau}_{\varepsilon}(\ell,\ell)\}_{\varepsilon>0,\ell\in\mathbb{N}}$, let $\tilde{\tau}_{0}(\infty,\infty)\coloneqq\lim_{\varepsilon\to 0}\lim_{\ell\to\infty}\tilde{\tau}_{\varepsilon}(\ell,\ell)$. Recall the set $\mathbf{H}$ in Corollary 4.2. Similarly to Theorem 5.2, we have the following convergence result.